OpenStudy (anonymous):

Explain how you would solve the equation -7 + x = 12 using the properties of equality

OpenStudy (anonymous):

x=19

OpenStudy (anonymous):

-7+x=12 add 7 to the LHS. Now, to balance the equality, add 7 to the right also. it becomes -7 + x +7 = 12 + 7 =>x+0=19 => x=19.

OpenStudy (anonymous):

add 7 to both sides

OpenStudy (unklerhaukus):

The property of equality that has been used is the property that say if you add some thing to both sides of an equality then you will still have equality

OpenStudy (anonymous):

@UnkleRhaukus exactly.!